School A and School B sent groups of students to a trip. School A rented and filled 2 mini buses and 3 coaches with 170 students. School B rented and filled 3 mini-buses and 2 coaches with 135 students. Every mini-buses had the same number of students in it, as did the coaches.
A) Write down two equations to represent the above information.
B) find the number of students in each mini-bus and in each coach

Answer by richwmiller(17219)

2m+3c=170,
3m+2c=135
9m+6c=405
4m+6c=340
5m=65
m=13
3m+2c=135
39+2c=135
2c=96
c=48